Opened 8 years ago

Last modified 7 years ago

#412 assigned flight processing

CR11/03, flight day 221/2011, Suert

Reported by: knpa Owned by: knpa
Priority: alpha 4 high Milestone: 2011 data processing completion
Component: Archiving Keywords:
Cc: Other processors:

Description (last modified by adbe)

Data location: /users/rsg/arsf/arsf_data/2011/flight_data/spain_portugal/CR11_03-2011_221_Suert/

Data arrived from ARSF via network transfer on 16/08/2011

CR11/03 Scientific objective:
1) to derive spatial information on vegetation, soil and terrain for hydrological model input by combined hyperspectral and LiDAR data
(2) to conduct a spectral fingerprinting approach allowing to trace back sources and pathways of sediments. Airborne data acquisition will be accompanied by an extensive field campaign.

CR11/03 Priority: a4h

PI: Hermann Kaufmann

Sensors:

Camera (16/11/2011)
Eagle (7/10/11)
Hawk (7/10/11)
LiDAR (7/10/11)

Change History (22)

comment:1 Changed 8 years ago by knpa

  • Description modified (diff)
  • Summary changed from CR11/03, flight day 221/2011, Campo to CR11/03, flight day 221/2011, Suert

comment:2 Changed 8 years ago by adbe

Processed Navigation

Now using alspp to process lidar

Generating hyperspectral config file

comment:3 Changed 8 years ago by adbe

LiDAR data is ready for delivery check, but there is no generated pdf logsheet yet

comment:4 Changed 8 years ago by adbe

Hyperspectral sct's found

Line sct
e1 -0.02
e2 -0.05
e3 -0.04
e4 -0.04
e5 -0.07
e6 -0.09
e7 -0.06
e8 -0.08
e9 -0.03
e10 -0.08
e11 -0.04
e12 -0.07
e13 -0.03
e14 -0.04
e15 -0.07
e16 -0.09
e17 -0.05
e18 -0.04
h1 0.65
h2 -0.08
h3 -0.04
h4 -0.06
h5 -0.04
h6 -0.06
h7 -0.03
h8 -0.06
h9 -0.04
h10 -0.06
h11 -0.04
h12 -0.02
h13 -0.04
h14 -0.05
h15 -0.04
h16 -0.06
h17 -0.02
h18 -0.05

comment:5 Changed 8 years ago by adbe

Hyperspectral data is ready for delivery check, but there is no generated pdf logsheet yet

comment:6 Changed 8 years ago by adbe

Created logsheet, LiDAR and hyperspectral is ready for delivery check

comment:7 Changed 8 years ago by emca

Starting Hyperspectral delivery check

comment:8 Changed 8 years ago by knpa

  • Owner set to adbe
  • Status changed from new to assigned

Started LiDAR delivery check.

comment:9 Changed 8 years ago by emca

Hyperspectral delivery check complete.

Filenames listed in readme do not match filenames in directory. Directories affected are fodis and level1b. Need to remove the '.bil_' from the filenames listed in the readme. e.g. e221xx1b.bil_mask.bil should be e221xx1b_mask.bil (create_latex_readme script has been fixed for this). Can you update or recreate the readme.

DEM: Viewed the DEM in ENVI - almost 1/3 contains no data values, the left side of the DEM. Maybe the 'west' value was too small on export from grass?
Also DEM name ('Suert.dem') in header file does not match the *.dem file name.

Otherwise, data all good.

comment:10 Changed 8 years ago by knpa

LIDAR delivery check completed.

Result: Fail.
I got to point 4 and determining that you need to fix the classification and stopped. Fix the other things too, let me know what you change and I'll recheck.

  1. Digital elevation section of DEM:

processing/delivery/CR11 03-221-lidar-20110831/dem/CR11 03-2011-221-lidar-wgs84_utm.dem and
processing/delivery/CR11 03-221-lidar-20110831/dem/CR11 03-2011-221-lidar-wgs84_latlong.dem
Both should have their baesname only.

  1. flightlines/ascii/LDR-CR11_03-2011-221-03.txt - these might need dos2txt (check with emma/mark)
  1. Should be more specific in data quality remarks about the problems with the first two lines. They are a third the size and look messed up because...
  1. CLASSIFICATION NOT GOOD ENOUGH. MOST OF THE TREES/BUSHES ARE FALSELY CLASSIFIED. LINES LOOKED AT: 5,6,7.

comment:11 Changed 8 years ago by adbe

  • I have loaded dem into grass and restricted the area to not include poor values
  • Changed name in header
  • Created ReadMe again

comment:12 Changed 8 years ago by adbe

  • I have removed processing/delivery/CR11 03-221-lidar-20110831/dem/ from ReadMe
  • I have done dos2txt on all txt files
  • I have been more explicit about the short lines
  • I have improved classification

comment:13 Changed 8 years ago by knpa

Second lidar delivery check:

Classification now acceptable.

No other problems found, ready for delivery.

comment:14 Changed 8 years ago by emca

Checked hyperspectral delivery again.

Readme is okay and DEM looks good now.

Ready to deliver.

comment:15 Changed 8 years ago by adbe

  • Description modified (diff)

LiDAR and Hyperspectral sent to Hermann Kaufmann on 7 October on disk 91.

comment:16 Changed 8 years ago by adbe

Camera ready for delivery check

comment:17 Changed 8 years ago by emca

Delivery checked Camera data - no problems. Ready to deliver.

comment:18 Changed 8 years ago by adbe

  • Description modified (diff)

Camera sent to Hermann Kaufmann on disk 119 on 16/11/2011

comment:19 Changed 8 years ago by mark1

Hyperspectral hawk bad pixel files and XML files have been re-created and re-delivered to PI (via email). This corrects sample number in bad pixel text files and Eagle IFOV value in XML and extra comments included.

comment:20 Changed 8 years ago by knpa

  • Component changed from Processing: general to Archiving
  • Owner changed from adbe to knpa

Preparing for archiving.

comment:21 Changed 8 years ago by knpa

Finished archiving preparation.

comment:22 Changed 7 years ago by knpa

Backed up by NEODC.

Moved to arsf archive disk.

Note: See TracTickets for help on using tickets.